# Diagnosing phase transitions through time scale entanglement

### v1

@@ Abstract
We propose a new method based on the compression of arbitrary multi-point correlators into quantics tensor trains and studying its so called bond dimension as an indicator of time scale entanglement in the system. This quantics tensor train diagnostics approach allows us to i) identify quantum phase transitions (ground state crossings) and thermal crossovers, ii) distinguish between these two by allowing flexibility through iii) using arbitrary available correlators and iv) the possibility to deal with low-accuracy input data through varying the quantics tensor train cutoffs. We verify the validity of the approach by studying the phase diagram of the Hubbard dimer, the four-site Hubbard ring with and without next-nearest neighbor hopping and the single-impurity Anderson model. We find time scale entanglement to be a system-inherent property becoming maximal at phase transitions and crossover, which translates to peaks in the quantics tensor train bond dimension of correlator functions. Additionally, the connection of the method to other physical quantities, such as various entanglement measures and susceptibilities, is shown.
@@ I Introduction
Phase transitions are among the most intriguing phenomena in physics. Yet, their numerical treatment, particularly in systems with nontrivial electronic behavior, can be challenging, especially when exotic phases emerge [1, 2, 3]. In particular, the full wave function is not available for many systems, and using susceptibilities to diagnose transitions can be brittle, since we need to know a priori where to look, i.e., we must make sure to compute the response associated with the transition in question or at least one that is not insensitive to the transition. In general, we can expect some but a much weaker signal in the susceptibilities that are not the one associated with the symmetry breaking for a second order phase transition. A first order phase transitions, on the other hand, will not be signaled at all in the susceptibilities prior to the phase transition.
Tensor trains, used originally as matrix product states (MPSs) for the ground state of spin systems [4, 5], have been elevated in the last years to a general compression ansatz [6, 7, 8]. By combining tensor trains with the quantics representation – a set of auxiliary qubits for exponentially different time or length scales – one obtains a compression ansatz for arbitrary functions in space and time [9, 10, 11] called a quantics tensor train (QTT). QTTs were instrumental, among others, in computational breakthroughs in modeling of turbulence [12, 13, 14, 15, 16], plasma physics [17], and quantum chemistry [18].
Some of the authors of this Letter have recently extended the use of QTTs for modeling electronic correlation functions [19], and found that it unifies prior pole [20, 21], basis [22, 23, 24, 25], and mesh refinement [26] ansätze into a single framework – one where doubling the size of the system amounts to adding a qubit, a mere linear increase. When combined with a way to sample QTTs [27, 28], this framework allows the solution of diagrammatic theories for systems which were previously inaccessible [29]. Moreover, full $n$-particle correlators can be compressed and stored.
The compression power of tensor trains depends on the connection strength between the qubits and thus between the length or time scales, encoded in the so-called bond dimension $D_{\mathrm{max}}$. For physical qubits used in MPSs for spin chains, this bond dimension has physical significance: it is related to the entanglement between two parts of the system [30]; for auxiliary qubits used in QTTs for correlation functions, it has until now only been used as a performance metric.
Figure 1: QTT bond dimension $D_{\mathrm{max}}$ of different imaginary-time correlation functions for the Hubbard dimer as a function of electronic repulsion $U$ and chemical potential $\mu/U$: (a) $D_{\mathrm{max}}$ for the one-particle Green’s function $G_{11}^{\uparrow\uparrow}$ at temperature $\beta^{-1}=1/50$, and (b,d) for the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ at $\beta^{-1}=1/50$ and $1/30$, respectively. For comparison, the value of the static, local spin susceptibility $\chi_{s}$ at $\beta^{-1}=1/50$ is shown in panel (c). Red dashed lines mark crossings of the ground state of the model. These crossings can be diagnosed as a sharp maximum in $D_{\mathrm{max}}$.
@@
In this Letter, we argue that the bond dimension of correlation functions has physical significance: it becomes maximal at crossovers and phase transitions [red dots in Fig. 1]. These maxima appear in both single- [Fig. 1 (a)] and two-particle propagators [Fig. 1 (b,c)], without the need to identify the specific susceptibility associated with the transition. We call this program of detecting phase transitions quantics tensor train diagnostics (QTTD).
@@ II Quantics tensor trains
In the QTT approach [6, 19] each variable is represented through a set of binary numbers or “quantics” corresponding to different length or time scales. The resulting tensor is then factorized into a tensor train (TT) at each scale. To illustrate the approach, let us consider a function $f(\tau)$ of a discretized variable $\tau$, with $\tau\in\{0,\ldots,M-1\}$ on $M=2^{R}$ grid points. In the quantics approach, $\tau$ is expressed in its binary representation $\tau=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},$ with $\sigma_{\ell}\in\{0,1\}\,$ leading to $f(\tau)$ being seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ instead. Now each tensor index $\sigma_{\ell}$ corresponds to an exponentially distinct length or time scale of the system. The first bit $\sigma_{1}$ represents the coarsest scale which divides the system into halves, while the last bit $\sigma_{R}$ reflects the finest scale.
@@
can be achieved by means of singular value decomposition (SVD) [19] or using the tensor cross interpolation (TCI) algorithm [8, 28, 27]. In Eq. (1), each $M_{\ell}$ is a $D_{\ell-1}\times 2\times D_{\ell}$ tensor with “physical” binary index $\sigma_{\ell}$ and virtual indices (“bonds”) $\alpha_{\ell-1},\alpha_{\ell}$, which are summed over. Eq. (1) translates to:
@@
in tensor network form [5]. We define the (maximum) bond dimension of the QTT as $D_{\mathrm{max}}=\textrm{max}_{\ell}(D_{\ell})$. Generally, the bond dimensions $D_{\ell}$ characterize the amount of “length or time scale entanglement” in the system are truncated in the TT factorization such that a specified error cutoff or tolerance $\epsilon$ is satisfied. In the case of SVD, usually the squared Frobenius norm $\epsilon=||F-\tilde{F}||^{2}_{\mathrm{F}}/||F||^{2}_{\mathrm{F}}$ is taken as a measure of the error of the approximation in Eq. (1). Many problems in many-body physics have been shown to use quantities that are strongly compressible with the QTT approach, which is possible if the bond dimension is sufficiently small [19, 31, 32, 33, 28, 29, 34, 35, 36, 37].
@@ III Quantics tensor trains diagnostics
Let us start from a general imaginary time two-point correlator $G_{AB}(\tau)=-\langle\hat{\mathcal{T}}\hat{A}(\tau)\hat{B}\rangle$ with the fermionic operators $\hat{A},\hat{B}$ and the time ordering operator $\hat{\mathcal{T}}$. Then, $G_{AB}(\tau)$ can be expressed in its Lehmann representation:
@@
where $1/\beta$ is temperature, $\tau\in[0,\beta)$ is imaginary time, $Z$ is the partition function, $\hat{H}$ is the Hamiltonian including the chemical potential, and $\hat{H}|n\rangle=\epsilon_{n}|n\rangle$ defines its eigenenergies and eigenstates. As can be seen, $G_{AB}(\tau)$ is the sum of exponentials. A single exponential has a QTT representation of bond dimension one, since it is a product of exponentials of the binary variables $\sigma_{l}$ at each scale: $e^{\tau\epsilon}=\prod_{\ell=1}^{R}e^{\epsilon\beta 2^{-\ell}\sigma_{\ell}}$. The bond dimension of the exact QTT representation of $G_{AB}$ is, thus, bound by the number of nonzero elements in the sum. Let us now consider the case of temperature approaching zero ($\beta\rightarrow\infty$), then
@@
First, let us focus on the case of a single non-degenerate ground state. Then, the bond dimension of the exact QTT representation is bound by the number of nonzero terms in the sum, so at maximum by $M-1$ (if all matrix elements are nonzero) for $M$ eigenstates $m$. Next, consider a ground state crossing of the two lowest states $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ ($N=2$). Then the bond dimension of the exact QTT is at most $22M-$ 2 since twice as many exponentials can contribute. In practice, it may be lower as only some of the states $|m\rangle$ will give nonzero matrix elements. Still, at the ground state crossing the sum consists of the exponentials and matrix elements associated with $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ leading to larger bond dimensions of $G_{AB}(\tau)$ than before or after the crossing. Therefore, at zero temperature, a quantum phase transition (QPT) or a ground state crossing of the Hamiltonian can be determined by monitoring the bond dimension of the QTT representation. We will call this approach quantics tensor train diagnostics (QTTD), where the diagnostics power of the approach is further examined in Appendix D. Similarly, ground state crossings of degenerate states (e.g. a doublet with a triplet state) will lead to maxima of the bond dimension in parameter space.
@@
Let us now consider finite, but low temperatures. Then, also excited states $|n\rangle$ contribute to $G_{AB}(\tau)$, suppressed by $e^{-\beta\epsilon_{n}}$. However, if we compress the two-point correlator into a QTT, most of these contributions will be truncated, if they are below a certain specified cutoff in the QTT construction. Hence, we will still see a peak of the maximum bond dimension in the vicinity of the corresponding QPT (at $T=0$), which will become less sharp with increasing temperature which smears out this transition.
@@
The QTTD approach can also be used to identify thermal crossovers of different regimes, such as the Kondo effect, that cannot be identified by merely looking at the lowest energy states in the ($T$-independent) eigenspectrum of the Hamiltonian. The reason for this is that at the crossover, both regimes with their excitations contribute, leading to larger bond dimensions. In contrast to the sharp peaks of the remnants of the QPTs (more precisely, ground state crossings), we expect the peaks in the QTT bond dimensions of the correlators at these thermal crossovers to be broader and their position to depend on temperature.
@@
Furthermore, since Fourier transformation to Matsubara frequencies can be represented by a low-rank matrix product operator acting on the QTT representation of the imaginary-time correlator [19, 38], it can be expected that peaks in the bond dimension of the frequency representation match those of their imaginary times counterparts. Since at a thermal crossover or at a QPT multiple phases or regimes provide diverse contributions to correlation functions, more information has to be exchanged between the different length or time scales of the system leading to an increase in length or time scale entanglement. Therefore, the QTTD approach is not limited to specific two-point correlators, but can be utilized with arbitrary nonzero multi-point correlators due to the QTT bond dimension as a measure of system-inherent time scale entanglement. Moreover, the peaks in the bond dimensions of the QTT representation of multi-point correlators become more pronounced and stable with respect to various QTT cutoffs than in the two-point case (see the derivation in App. A). In App. B we also elaborate on how to treat low-accuracy data of correlators by varying the QTT cutoff to more precisely identify phase transitions and crossovers seen as stable peaks in $D_{\mathrm{max}}$.
@@
We can summarize the QTTD procedure in the following way. i) Calculate an available correlator of the system. ii) Vary the QTT cutoffs to find stable peaks corresponding to possible phase transitions or crossovers. iii) Iterate on that procedure for different available correlators to verify the obtained maxima as universal features corresponding to the system-inherent rise in time scale entanglement associated with a phase transition or crossover. iv) Vary the temperature in the calculation of the correlators to disentangle ground state crossings from thermal crossovers. Moreover, since various physical properties, e.g. physical susceptibilities or entanglement, heavily depend on the dominant phase, features of these properties can also be visible in the maxima in the QTT bond dimension of the correlators.
@@ IV Finite one-dimensional Hubbard models
Let us illustrate QTTD by attempting to find ground state crossings of a series of small benchmark systems. We start with the one dimensional Hubbard model on $N$ sites:
$\hat{H}$ $=t\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+1,\sigma}+\hat{c}_{i+1,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)+t^{\prime}\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+2,\sigma}+\hat{c}_{i+2,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)$
$+\sum_{i=1}^{N}\left(U\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\mu(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})\right),$ (5)
with the (next-)nearest neighbor hopping amplitude $t(t^{\prime})$, the local on-site Coulomb interaction $U$, the chemical potential $\mu$, the number operators $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ defined via the fermionic annihilation (creation) operators $\hat{c}_{i,\sigma}^{(\dagger)}$ with $\sigma=\uparrow,\downarrow$ and site index $i$. We use $t\equiv k_{B}\equiv 1$ to set energy and temperature units. In the following, we will study the cases of the Hubbard dimer and the four-site Hubbard ring with and without next-nearest neighbor hopping.

### v2

@@ Abstract
Spatial entanglement of wave functions has matured into an enthralling and very active research area. Here, we unearth a completely different kind of entanglement, the entanglement between different time scales. This is feasible through quantics tensor train diagnostics (QTTD), wherein the bond dimension for an $n$-particle correlation function allows diagnosing the temporal entanglement. As examples, we study time-scale entanglement of the Hubbard dimer, the four-site Hubbard ring with and without next-nearest neighbor hopping and the single-impurity Anderson model. Besides introducing the QTTD method, our major finding is that the time-scale entanglement is generically maximal at phase transitions and crossovers. This is independent of the correlation function studied. Thus, QTTD is a universal tool for detecting quantum phase transitions, ground state crossings in finite systems, and thermal crossovers.
@@ I Introduction
Phase transitions are among the most intriguing phenomena in physics. Yet, their numerical treatment and detection can be challenging, especially when exotic phases emerge [1, 2, 3]. In extended many-body systems, the full wave function is usually not available, and using susceptibilities to diagnose transitions can be brittle, since we need to know a priori where to look, i.e., we must make sure to compute the response associated with the transition in question or at least one that is not insensitive to the transition. In general, we can expect some but a much weaker signal in the susceptibilities that are not the one associated with the symmetry breaking for a second order phase transition. A first order phase transition, on the other hand, will not be signaled at all in the susceptibilities prior to the phase transition.
Similarly, entanglement is one of the most enigmatic quantum phenomena. It is equally challenging to calculate in true many-electron systems where many electrons contribute and are entangled [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. In (quasi-) one-dimensional spin systems, matrix product states (MPS) [14, 15] have unlocked much of the theoretical and numerical analysis of entanglement [16, 17] and its growth [18, 19, 20]. MPS add to the physical space an auxiliary, latent space, whose dimension (“bond dimension”) indicates the strength of spatial entanglement [21] [22].
Recently, it was shown that the machinery of MPS can be reused to compactify arbitrary functions in space and time [23, 24, 25, 26, 27, 28]. Instead of a physical spin at some site, each tensor represents a different time/length scale of the function, so as we step along the train, we zoom in and out rather than move left or right. This ansatz was termed quantics tensor train (QTT). It was instrumental in breaking computational barriers in modeling turbulence [29, 30, 31, 32, 33], plasmas [34], quantum chemistry [35], and – most relevant to this Letter – electronic correlation functions. In this latter case, some of us showed how to sample [36, 37], store [38], and perform calculations [39] with electronic response functions which were previously inaccessible. QTTs by construction also feature a latent space and an associated bond dimension, which has hitherto been chiefly used as a measure of performance.
Figure 1: Entanglement of exponentially different time scales for the Hubbard dimer at $T=0.02$: (a) Local spin susceptibility $\chi_{s}$ indicating different ground states filled with one, two and three electrons; (b) QTT maximum bond dimension $D_{\max}:=\max_{\ell}D_{\ell}$ of the imaginary-time four-point Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ as a function of interaction strength $U$ and chemical potential $\mu/U$, where red-dashed lines mark crossings of the ground state; and (c) QTT bond dimension $D_{\ell}$ between timescale $2^{-\ell}$ and $2^{-\ell-1}$ for different values of $\mu/U$ at $U=8$, marked by the same symbol and color in (a) and (b).
In this Letter, we reason by analogy with MPS, where the bond dimension gauges spatial entanglement: we propose the QTT bond dimension of electronic functions as a physical measure, namely as the strength of time and length scale entanglement. For the example of the Hubbard dimer, we plot this over different time scales in Fig. 1(c).
@@
Furthermore, we observe that the maximal bond dimension $D_{\mathrm{max}}$ shows a maximum at crossovers and phase transitions [red dots in Fig. 1(b)]. These maxima appear in both single- and two-particle propagators [Fig. 1 (b)], without the need to identify the specific susceptibility associated with the transition. We call this program of detecting phase transitions quantics tensor train diagnostics (QTTD).
@@ II Quantics tensor trains
In QTTs  [23, 38] each variable is represented through a set of binary numbers or “quantics” corresponding to different length or time scales. The resulting tensor is then factorized into a tensor train (TT) at each scale. For illustration, consider a function $f(\tau)$ of a discretized variable $\tau$, with $\tau\in\{0,\ldots,M-1\}$ on $M=2^{R}$ grid points. In quantics, $\tau$ is expressed in binary representation $\tau=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},$ with $\sigma_{\ell}\in\{0,1\}\,$ leading to $f(\tau)$ being seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ instead. Now each tensor index $\sigma_{\ell}$ corresponds to an exponentially distinct length or time scale of the system. The first bit $\sigma_{1}$ represents the coarsest scale which divides the system into halves, while the last bit $\sigma_{R}$ reflects the finest scale.
@@
can be achieved with singular value decomposition (SVD) [38] or using tensor cross interpolation (TCI) [25, 37, 36]. In Eq. (1), each $M_{\ell}$ is a $D_{\ell-1}\times 2\times D_{\ell}$ tensor with “physical” binary index $\sigma_{\ell}$ and virtual indices (“bonds”) $\alpha_{\ell-1},\alpha_{\ell}$, which are summed over. Eq. (1) translates to:
@@
[…]
@@ III Quantics tensor trains diagnostics
Let us start from a general imaginary time two-point correlator $G_{AB}(\tau)=-\langle\hat{\mathcal{T}}\hat{A}(\tau)\hat{B}\rangle$ with fermionic operators $\hat{A},\hat{B}$ and the time ordering operator $\hat{\mathcal{T}}$. Then, $G_{AB}(\tau)$ can be expressed in its Lehmann representation:
@@
where $1/\beta$ is temperature, $\tau\in[0,\beta)$ is imaginary time, $Z$ is the partition function, $\hat{H}$ is the Hamiltonian including the chemical potential, and $\hat{H}|n\rangle=\epsilon_{n}|n\rangle$ defines its eigenenergies and eigenstates. Clearly, $G_{AB}(\tau)$ is the sum of exponentials. An exponential has a QTT representation of bond dimension one, since it is a product of exponentials of the binary variables $\sigma_{l}$ at each scale: $e^{\tau\epsilon}=\prod_{\ell=1}^{R}e^{\epsilon\beta 2^{-\ell}\sigma_{\ell}}$. The bond dimension of the exact QTT representation of $G_{AB}$ is, thus, bound by the number of nonzero elements in the sum. Let us now consider the case of temperature approaching zero ($\beta\rightarrow\infty$), then
@@
For a single non-degenerate ground state, the bond dimension of the exact QTT representation is bounded by the number of nonzero terms in the sum, maximally by M1M-1 for $M$ eigenstates $m$. Next, consider a ground state crossing of the two lowest states $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ ($N=2$). Then the bond dimension of the exact QTT is at most 2M22M-2 since twice as many exponentials can contribute. In practice, it may be lower as fewer states $|m\rangle$ will give nonzero matrix elements. Still, at the ground state crossing the sum consists of the exponentials and matrix elements associated with $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ leading to larger bond dimensions of $G_{AB}(\tau)$ than before or after the crossing. Therefore, at zero temperature, a quantum phase transition (QPT) or a ground state crossing can be determined by monitoring the bond dimension of the QTT representation. We call this approach quantics tensor train diagnostics (QTTD), where the diagnostics power of the approach is further examined in Appendix D. Similarly, ground state crossings of degenerate states (e.g. a doublet with a triplet state) will lead to maxima of $D_{\mathrm{max}}$ in parameter space.
@@
Consider finite, but low temperatures. Then, also excited states $|n\rangle$ contribute to $G_{AB}(\tau)$, suppressed by $e^{-\beta\epsilon_{n}}$. Most of these contributions will be truncated, if they are below a certain specified cutoff in the QTT construction of the two-point correlator. Hence, we will still see a peak of $D_{\mathrm{max}}$ in the vicinity of the corresponding QPT (at $T=0$), that will be smeared out with increasing temperature.
@@
QTTD can also identify thermal crossovers of different regimes, such as the Kondo effect, that cannot be identified by merely looking at the lowest energy states in the ($T$-independent) eigenspectrum of the Hamiltonian. At the crossover, both regimes with their excitations contribute, leading to larger bond dimensions. In contrast to the sharp peaks of the remnants of QPTs (more precisely, ground state crossings), we expect the peaks of $D_{\mathrm{max}}$ of the correlators at thermal crossovers to be broader and their position to depend on temperature.
@@
Since at thermal crossovers or at QPTs multiple phases or regimes provide diverse contributions to correlation functions, more information has to be exchanged between different length or time scales of the system leading to an increase in length or time scale entanglement. Therefore, QTTD is not limited to specific two-point correlators, but can be utilized with arbitrary nonzero multi-point correlators due to the QTT bond dimension as a measure of system-inherent time scale entanglement. Moreover, peaks in $D_{\mathrm{max}}$ of the QTT representation of multi-point correlators become more pronounced and stable with respect to various QTT cutoffs than in the two-point case (see derivation in App. A). In App. B, we also elaborate on how to treat low-accuracy data of correlators by varying the QTT cutoff to more precisely identify phase transitions and crossovers seen as stable peaks in $D_{\mathrm{max}}$.
@@
We can summarize the QTTD procedure in the following way. i) Calculate an available correlator of the system. ii) Vary the QTT cutoffs to find stable peaks corresponding to possible phase transitions or crossovers. iii) Iterate on this procedure for different available correlators to verify the obtained maxima as universal features corresponding to the system-inherent rise in time scale entanglement associated with a phase transition or crossover. iv) Vary the temperature in the calculation of the correlators to disentangle ground state crossings from thermal crossovers. Moreover, since various physical properties, e.g. physical susceptibilities or entanglement, heavily depend on the dominant phase, features of these properties can also be visible in the maxima in the QTT bond dimension of the correlators.
@@ IV Finite one-dimensional Hubbard models
We start with the one dimensional Hubbard model on $N$ sites:
$\hat{H}=-\sum_{\sigma}\sum_{i,j=1}^{N}t_{i-j}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}+\sum_{i=1}^{N}\left(U\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\mu(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})\right),$ (5)
where $t_{\pm 1}\equiv t$ and $t_{\pm 2}\equiv t^{\prime}$ are the nearest and next-nearest hopping amplitude, respectively, all other hopping amplitudes are set to zero, the local on-site Coulomb interaction is $U$, the chemical potential is $\mu$, and the number operators $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ are defined via the fermionic annihilation (creation) operators $\hat{c}_{i,\sigma}^{(\dagger)}$ with $\sigma=\uparrow,\downarrow$ and site index $i$. We use $t\equiv k_{B}\equiv 1$ to set energy and temperature units. In the following, we will study the cases of the Hubbard dimer and the four-site Hubbard ring with and without next-nearest neighbor hopping.